# Surgical Text-to-Image Generation

### v1

@@ Abstract
Acquiring surgical data for research and development is significantly hindered by high annotation costs and practical and ethical constraints. Utilizing synthetically generated images could offer a valuable alternative. In this work, we conduct an in-depth analysis on adapting text-to-image generative models for the surgical domain, leveraging the CholecT50 dataset, which provides surgical images annotated with surgical action triplets (instrument, verb, target). We investigate various language models and find T5 to offer more distinct features for differentiating surgical actions based on triplet-based textual inputs. Our analysis demonstrates strong alignment between long and triplet-based captions, supporting the use of triplet-based labels. We address the challenges in training text-to-image models on triplet-based captions without additional input signals by uncovering that triplet text embeddings are instrument-centric in the latent space and then, by designing an instrument-based class balancing technique to counteract the imbalance and skewness in the surgical data, improving training convergence. Extending Imagen, a diffusion-based generative model, we develop Surgical Imagen to generate photorealistic and activity-aligned surgical images from triplet-based textual prompts. We evaluate our model using diverse metrics, including human expert surveys and automated methods like FID and CLIP scores. We assess the model performance on key aspects: quality, alignment, reasoning, knowledge, and robustness, demonstrating the effectiveness of our approach in providing a realistic alternative to real data collection.
@@ 4.1 Experiment
The 321.9321.9{321.9}321.9M parameters Surgical Imagen is trained for 300K iterations with a batch size of $16$, using data parallelism to leverage 4×4$\times$4 ×V100 GPUs (64GB each). Our training set included approximately 101K image-text pairs from the CholecT50 dataset [1]. The initial output dimensions were $64\times 64\times 3$, with super-resolution scaling by a factor of 4. We used Adafactor for the base model and Adam for the super-resolution model as in [7]. Inference involved 1000 diffusion steps to generate the final high-quality images.
@@ 4.2 Evaluation
For human evaluation, we design a study in form of an online survey polling surgeons and medical professionals, with varying degree of expertise and levels of experience in laparoscopic cholecystectomy, to assessed the photorealism and alignment of the generated images with the text prompts. The survey comprised 140 questions, each accompanied by four image options, totaling 640 images — approximately evenly split between real and generated. The survey was divided into two phases, taking around 9 minutes to complete. In the first phase, evaluators were asked to identify real images from a mix of actual and generated images across 5 randomly selected questions. In the second phase, consisting of 10 randomly selected questions, they were tasked with selecting images that best matched given prompts, focusing on assessing the alignment of tools, actions of tools, anatomies, and tool-tissue interactions. Examples of survey questions include: ”- Which of these images represents a coagulation of the liver? - Select the images corresponding to a grasper retracting the gallbladder. - Select the images where a tool is used to dissect the cystic plate. - Choose only the real images from options below. We analyzed the percentage of real and generated images identified as photorealistic and correctly aligned by respondents to gauge the model’s performance effectively. The survey and the responses can be accessed via the project page.
@@ 4.3.1 Quality
Quality evaluates the photo-realism or visual realism of the generated surgical images to ensure they closely resemble authentic images, crucial for applications requiring accurate and convincing representations of surgical scenarios. Here we employ both human expert and automated evaluation. For the human evaluation, an initial survey with 13 respondents is analyzed in Table 1. Surgical Imagen achieves $34.6\%$ preference rate indicating high image quality generation compared to $57.7\%$ preference to the actual images. This shows how closely the generated images are to the real image such that $42.3\%$ of the time, experts fails to identify the real images and $34.6\%$ of the time, they classifies the synthetic images as real.
For the automated evaluation, we report the FID score in Table 2 showing Surgical Imagen on $3.37$ compare to baselines.
@@
Real $57.7\pm 0.20$ $72.3\pm 0.31$
Surgical Imagen $34.6\pm 0.47$ $43.6\pm 0.71$
@@
Real - - - $23.01\pm 0.11$
StackGAN [37] $5.83$ $28.53\pm 0.86$
Surgical Imagen $3.70$ $26.84\pm 0.45$
@@ 4.3.2 Task Alignment
This assessment is aim at verifying the semantic coherence between the generated images and the accompanying input text prompts, ensuring that the visual content accurately corresponds to the intended surgical context for effective communication and understanding. We also employ both human expert and automated evaluations to measure the degree of text-image alignment in the generated images. For the expert evaluation, we measure how often the clinicians think that the AI generated images are aligned with the input text semantics. As reported in Table 1, the clinicians agree with Surgical Imagen $43.6\%$ of the time, meaning that the rate of disagreement is also similar ($56.4$) suggesting our model’s ability to generate images that align well with real-world surgical activity based on three words captions. On the other hand, the clinicians rate the real images alignment as correct $72.3\%$ of the time, leaving room for doubt ($27.7\%$) in such complex domain as surgery.
@@
Overall, the medical participants were highly challenged by the realistic characteristics of the samples, which validates their authenticity. The survey in Table 1 will be updated with more responses from the wider endoscopist community.
@@ 5.1.1 Simulation
Simulation stands as the primary domain in which we anticipate the widespread adoption of Surgical Imagen. A pivotal forthcoming advancement of the model would empower us to generate complete videos, thus sparking discourse within the simulation community. These materials are poised to play a pivotal role in broadening access to an extensive array of synthetic surgical laparoscopic procedures, thereby democratizing their availability. The capability to produce images affords us the opportunity to replicate frames depicting laparoscopic complications, a facet often underrepresented in existing datasets due to data scarcity. Beyond addressing complications, our system excels in manipulating camera angles, thereby reproducing sub-optimal views: a feature that has already demonstrated its efficacy in reducing the learning curve within surgical training environments [41]. Additionally, we are exploring future avenues that extend simulation into the realm of virtual reality (VR). This entails training and refining the model using publicly accessible stereo view datasets sourced from DaVinci surgeries, with the objective of generating paired images. These images could subsequently be integrated into VR headsets for immersive 3D simulation, promising enhanced training efficacy characterized by reduced time investment and errors [42].
@@ Acknowledgements
The authors acknowledge the following clinicians for their participation in filling the evaluation survey: Alain Garcia, Florent Alexandre, Giovanni Gulglielmo Laracca, Giuseppe Massimiani, Juan Manuel Verde, Julieta Montanelli, Marta Goglia, Matteo Pavone, Leonardo Sosa Valencia.

### v2

@@ Abstract
Acquiring surgical data for research and development is significantly hindered by high annotation costs and practical and ethical constraints. Utilizing synthetically generated images could offer a valuable alternative. In this work, we explore adapting text-to-image generative models for the surgical domain using the CholecT50 dataset, which provides surgical images annotated with action triplets (instrument, verb, target). We investigate several language models and find T5 to offer more distinct features for differentiating surgical actions on triplet-based textual inputs, and showcasing stronger alignment between long and triplet-based captions. To address challenges in training text-to-image models solely on triplet-based captions without additional inputs and supervisory signals, we discover that triplet text embeddings are instrument-centric in the latent space. Leveraging this insight, we design an instrument-based class balancing technique to counteract data imbalance and skewness, improving training convergence. Extending Imagen, a diffusion-based generative model, we develop Surgical Imagen to generate photorealistic and activity-aligned surgical images from triplet-based textual prompts. We assess the model on quality, alignment, reasoning, and knowledge, achieving FID and CLIP scores of 3.7 and 26.8% respectively. Human expert survey shows that participants were highly challenged by the realistic characteristics of the generated samples, demonstrating Surgical Imagen’s effectiveness as a practical alternative to real data collection.
footnotetext: Under consideration at Pattern Recognition Letters
@@ 4.1 Experiment
The 321.9321.9{321.9}321.9M parameters Surgical Imagen is trained for 300K iterations with a batch size of $16$, using data parallelism to leverage 4×4$\times$4 ×V100 GPUs (32GB each). Our training set included approximately 101K image-text pairs from the CholecT50 dataset [1]. The initial output dimensions were $64\times 64\times 3$, with super-resolution scaling by a factor of 4. We used Adafactor for the base model and Adam for the super-resolution model as in [7]. Inference involved 1000 diffusion steps to generate the final high-quality images.
@@ 4.2 Evaluation
For human evaluation, we design a study in form of an online survey polling surgeons and medical professionals, with varying degree of expertise and levels of experience in laparoscopic cholecystectomy, to assessed the photorealism and alignment of the generated images with the text prompts. The survey comprised 140 questions, each accompanied by four image options, totaling 640 images — approximately evenly split between real and generated. The survey was divided into two phases, taking around 9 minutes to complete. In the first phase, evaluators were asked to identify real images from a mix of actual and generated images across 6 randomly selected questions. In the second phase, consisting of 10 randomly selected questions, they were tasked with selecting images that best matched given prompts, focusing on assessing the alignment of tools, actions of tools, anatomies, and tool-tissue interactions. Examples of survey questions include: ”- Which of these images represents a coagulation of the liver? - Select the images corresponding to a grasper retracting the gallbladder. - Select the images where a tool is used to dissect the cystic plate. - Choose only the real images from options below. We analyzed the percentage of real and generated images identified as photorealistic and correctly aligned by respondents to gauge the model’s performance effectively. The survey and the responses can be accessed via the project page.
@@ 4.3.1 Quality
Quality evaluates the photo-realism or visual realism of the generated surgical images to ensure they closely resemble authentic images, crucial for applications requiring accurate and convincing representations of surgical scenarios. Here we employ both human expert and automated evaluation. For the human evaluation, an initial survey with 60 respondents is analyzed in Table 1. Surgical Imagen achieves $39.0\%$ preference rate indicating high image quality generation compared to $66.3\%$ preference to the actual images. This shows how closely the generated images are to the real image such that $33.7\%$ of the time, experts fails to identify the real images and $39.0\%$ of the time, they classifies the synthetic images as real.
For the automated evaluation, we report the FID score in Table 2 showing Surgical Imagen on $3.7$ compare to baselines.
@@
Real $66.3\pm 0.2$ $69.3\pm 0.2$
Surgical Imagen $39.0\pm 0.1$ $51.5\pm 0.2$
@@
Real - - - $23.01\pm 0.1$
StackGAN [37] $5.83$ $28.53\pm 0.9$
Surgical Imagen $3.70$ $26.84\pm 0.5$
@@ 4.3.2 Task Alignment
This assessment is aim at verifying the semantic coherence between the generated images and the accompanying input text prompts, ensuring that the visual content accurately corresponds to the intended surgical context for effective communication and understanding. We also employ both human expert and automated evaluations to measure the degree of text-image alignment in the generated images. For the expert evaluation, we measure how often the clinicians think that the AI generated images are aligned with the input text semantics. As reported in Table 1, the clinicians agree with Surgical Imagen $51.5\%$ of the time, meaning that the rate of disagreement is even less ($48.5$) suggesting our model’s ability to generate images that align well with real-world surgical activity based on three words captions. On the other hand, the clinicians rate the real images alignment as correct $69.3\%$ of the time, leaving room for doubt ($30.7\%$) in such complex domain as surgery.
@@
Overall, the medical participants were highly challenged by the realistic characteristics of the samples, which validates their authenticity. The survey is left open for more responses from the global endoscopist community.
@@ 5.1.1 Simulation
Simulation stands as the primary domain in which we anticipate the widespread adoption of Surgical Imagen. A pivotal forthcoming advancement of the model would empower us to generate complete videos, thus sparking discourse within the simulation community. These materials are poised to play a pivotal role in broadening access to an extensive array of synthetic surgical laparoscopic procedures, thereby democratizing their availability. The capability to produce images affords us the opportunity to replicate frames depicting laparoscopic complications, a facet often underrepresented in existing datasets due to data scarcity. Beyond addressing complications, our system excels in manipulating camera angles, thereby reproducing sub-optimal views: a feature that has already demonstrated its efficacy in reducing the learning curve within surgical training environments [41]. Furthermore, the application of Surgical Imagen extends to virtual reality (VR). This entails training and refining the model using publicly accessible stereo view datasets sourced from robotic surgeries (e.g. DaVinci), with the objective of generating paired images. These images could subsequently be integrated into VR headsets for immersive 3D simulation, promising enhanced training efficacy characterized by reduced time investment and errors [42].
@@ Acknowledgements
The authors acknowledge the following clinicians for their participation in filling the evaluation survey: Adolfo Gutierrez, Alain Garcia (IHU Strasbourg), Alicia López (A Coruña University Hospital), Amir Ashraf Ganjouei (UCSF), Ana (Otero), Andrea Balla (University Hospital Virgen Macarena), Andrea Rosati (Fondazione Policlinico Gemelli, Rome), Anna Minasyan (Coruña University Hospital, Coruña, Spain), AS Soares (Hospital prof. Dr. Fernando Fonseca, UCL), Clarisa (Ponderas Academic Hospital), Donald Van Der Fraenen (Maria middelares Gent), Federica (University of Turin), Felipe Asenjo de Castro (Hospital Central de la Defensa “Gomez Ulla” Madrid) Florent Alexandre (Strasbourg IHU), Francesco Marzola (University of Turin - MITIC), Gabriel Szydlo Shein (Université de Strasbourg), Georgios Kourounis (Newcastle University), Giovanni Guglielmo Laracca (Sant’Andrea Hospital, Sapienza University of Rome, Italy), Giuseppe Massimiani (Università Cattolica del Sacro Cuore), Helmuth Radrich (Karl Storz VentureONE), Joris Vangeneugden (Ghent University Hospital), Juan Manuel Verde (IHU Strasbourg), Julieta Montanelli (IHU Strasbourg), Lachlan Dick (University of Edinburgh), Leonardo Sosa Valencia, Luca Gordini (ASL Sulcis), Luigi Boni (Fondazione IRCCS Ca’ Granda Ospedale Maggiore Policlinico di Milano, University of Milan), Manish Jethani (Kokilaben Dhirubhai Ambani Hospital Mumbai India), Marta Goglia (Sapienza University of Rome), Matteo Pavone (Policlinico Gemelli/IHU Starsbourg/IRCAD Strasbourg), Michele Ammendola (’Magna Graecia’ University & ’Renato Dulbecco’ University Hospital, Catanzaro, Italy), Muhammad Umar Younis (Mediclinic City Hospital Dubai), Nabani Banik (Université de Strasbourg), Patricia Deltell (Hospital general Alicante), Pietro Leoncini (University of Turin MITIC), Po-Hsing Chiang (Chang Gung Memorial Hospital, Taiwan), Roberto Spagnulo (MITIC- UniTO), Romina (Braverman), Salvador Morales-Conde (University Hospital Virgen Macarena, Sevilla, Spain), Sofia Teixeira da Cunha (Kantonsspital Olten), Sonia Pérez-Bertólez (Hospital Sant Joan de Déu), Yanhua Zhang (PoliTo) and others.